Skip to end of metadata
Go to start of metadata

s4, s8, s10, s15 all coming in.  s17 had a few samples come through Sat afternoon, but nothing recent.

s8 IP address: 166.255.144.44

Ott: all okay, reporting 0.

TRH: s8 TRH is reporting -99 for Rfan, which is supposed to mean that the fan is not turning.  Others okay.

P: all okay

CSAT: all okay

EC150: s8 all bad values (including Tirga and Pirga)

Victron: all okay (need to look at parsed DIDs 61-65 to get all fields)

Gill 2D: all okay

NR01: all okay

Tsoil: all okay

Gsoil: all okay (all same sign, so far)

Qsoil: all okay

TP01: all okay

** Need boom directions for CSATs and Gills to process directions properly


1 Comment

  1. Steven Oncley AUTHOR

    This morning, Kurt said that s17 should have been set up.  In the raw data, I see values from 00:19-00:39 UTC (on Mar 8).  Looking at these 20 min of data:

    Ott: okay

    TRH: okay

    P: not reporting (presumably the 7E1 issue)

    CSAT/EC150: okay

    Victron: not reporting

    Gill 2D: okay

    NR01: okay

    4 soils: okay

    dyndns is reporting the last IP address was 166.255.144.45, but not pinging now.